OpenStudy (anonymous):

4.What is the sum of an 8-term geometric sequence if the first term is 10 and the last term is 781,250? i need help findng the common ratio ? please

OpenStudy (anonymous):

8th term = ar^7 and first term = a r^7 = = ar^7 / a = r^7 = 781250 / 10 = 78125 r = seventh root of 78125 = 5 common ratio = 5

OpenStudy (anonymous):

can you calculate the sum?

OpenStudy (anonymous):

yes i just needed help finding the common ratio, thank you soo much

OpenStudy (anonymous):

yw